# PosePilot: Steering Camera Pose for Generative World Models with Self-supervised Depth

### v1

@@ Abstract
world model, video generation, autonomous driving
@@ I Introduction
In recent years, great advancements have been made in the development of autonomous driving systems. While existing AD methods [1, 2] demonstrate promising results in various scenarios, they still face significant challenges when confronted with long-tail distribution or out-of-distribution situations. These edge cases, which occur infrequently in typical training datasets but are critical in real-world driving environments, expose limitations in the robustness and generalization ability of current models. One promising solution is world models [3, 4, 5, 6, 7, 8, 9, 10], which capture the structure and dynamics of an environment, enabling an autonomous driving system to simulate and predict future states by ”imagining” the external world. Through building such models, these generative models can help AD systems anticipate future states, reason about complex dynamics and better generalize to novel or unexpected situations.
@@
  • We introduce PosePilot, a light-weight camera pose control method that enhances the pose controllability in world models, which can be easily applied to other world models to generate videos with flexible camera viewpoints.
  • We propose a photometric loss, which integrates self-supervised depth and ego-motion estimation to enhance the structural and motion understanding of world models.
  • We conduct comprehensive experiments and ablation study to analyze PosePilot, which boasts the controllability for both diffusion-based and autoregressive-based generative world models on autonomous driving and real world datasets.
@@ II-A World Models for Driving Scenes
World Models [20, 21] are generative models that predict future frames to simulate dynamics. Recent advances enabling video generation to augment autonomous driving data with diverse road conditions and rare corner cases [22, 23]. Existing methods use various conditions including images, text, 3D layouts, and actions to generate specific driving scenes or predict future video sequences. GAIA-1 [24] generates realistic and diverse driving videos by integrating video, text, and action inputs, demonstrating a strong understanding of contextual information and physical principles. DriveDreamer [25] improves driving scenario generation by adding HD maps and 3D boxes for better video quality and enabling the generation of future driving actions, aiding in decision-making. WorldDreamer [26] frames world modeling as an unsupervised challenge within visual sequence modeling, inspired by large language models. MUVO [27] enhances world modeling by integrating LIDAR point clouds, improving the prediction of driving environments and generating 3D occupancy grids that integrate well with downstream tasks. OccWorld [28] leverages 3D occupancy data to predict environmental evolution and guide autonomous vehicle actions, marking a shift toward multimodal approaches in autonomous driving research. In contrast, our work does not design a brand new framework for world model but serves as a plug-and-play module that builds upon existing world models to provide effective camera pose control during video generation for driving scenes.
@@ III Methodology
We propose PosePilot, a lightweight camera control enhancer in generative world models, focusing on enhancing the camera controllability in off-the-shelf world models for autonomous driving. As illustrated in Fig. 2, PosePilot is comprised of three losses: two photometric control losses and a pose regression loss. We adapt PosePilot  on top of the powerful generative world models [34, 10, 35].
III-A Photometric Control Loss
In contrast to traditional video generation models, a key factor in the world models for autonomous driving is the camera parameters. Once the extrinsic matrices (or pose) of each camera are available, we can use the structure-from-motion methods to generate the structural and dynamic context to guide the video generation. Inspired by monocular self-supervised depth estimation [17], we incorporate a photometric control loss into our generative model. The objective of the photometric control loss is to derive real-world scale information from the camera’s extrinsic matrices, which provides prior information for video generation.
@@ III-A Photometric Control Loss
III-B Inverse Photometric Control Loss
Similar to the photometric control loss in the previous subsection, we introduce an inverse photometric control loss to further improve the video-pose consistency of the generated frames. Instead of solely warping the original frame to the viewpoint of the generated frame, we also inverse warp the generated frame back to the original viewpoint. This bidirectional alignment helps reduce the artifacts due to occlusions and dis-occlusions.
Let us denote the generated frames as $(\hat{f}^{i},\hat{f}^{j})$, which are obtained from the world model. The learnable depth readout provides their respective depth maps $(D^{i},D^{j})$, and the pose readout outputs the relative 6D camera transformations $T^{i\rightarrow j}$ and $T^{j\rightarrow i}$. To map $\hat{f}^{j}$ back to $\hat{f}^{i}$, we warp using the same bilinear sampling mechanism as above. Specifically, the pixel projection from $\hat{f}^{j}$ to the warped inverse view $\hat{f}^{j\rightarrow i}$ follows the same procedure as in the forward warping. Denoting $x^{j}$ as a pixel in the generated frame $\hat{f}^{j}$, its inverse-warped coordinate $x^{j\rightarrow i}$ is found by:
@@ III-B Inverse Photometric Control Loss
To ensure the pose readouts generate precise camera pose, we also utilize a regression loss $L_{MSE}$ to refine the pose readouts.
III-C Total Loss
We combine the forward and inverse photometric control losses with our generative objective into a comprehensive total loss. Let $L_{g}$ represent the generative objective (e.g., the diffusion-based reconstruction or adversarial loss), $L_{p}$ be the forward photometric control loss, and $L_{p}^{\text{inv}}$ be the inverse photometric control loss. We employ weighting factors $\alpha_{p}$ and $\alpha_{p^{\text{inv}}}$ to balance the relative importance of these terms, which yields the total loss:
@@ III-C Total Loss
where the $L_{MSE}$ represents the regression loss to the pose readouts and $\alpha_{mse}$ is its factor, explicitly refining the pose estimation to achieve more precise pose control. In practice, $\alpha_{p}$, $\alpha_{p^{\text{inv}}}$ and $\alpha_{mse}$ can be tuned to find an optimal trade-off between the generative fidelity of the generative model and the geometric consistency enforced by both the forward and inverse photometric control terms. By minimizing $L_{\text{total}}$, the framework simultaneously learns to produce high-quality, realistic frames while ensuring that the predicted depth and pose remain consistent in both directions of warping.
@@ IV-B Main Results
We adapt our methods to the existing world model solutions on the nuScenes validation set. We divide the nuScenes validation set into four subsets. The TranErr and RotErr scores are measured on each subset and then averaged. We conduct experiments on most of the publicly available world models, including diffusion-based methods [34, 10] and auto-regression-based approach [35]. The results are shown in Tab. I. We can see that with our PosePilot, all of the models can get relatively lower TranErr & RotErr, which demonstrates the effectiveness of the proposed architecture. For example, in diffusion-based methods, with our PosePilot, Vista gets a 0.31 decrease on translation error and 0.21 decrease on rotation error. In auto-regression-based approach, Drivingworld also gets a 0.22 decrease on translation error and 0.16 decrease on rotation error, indicating the superior generalization of our method. We also show the qualitative results of PosePilot at Fig. 3, we can see that our PosePilot  contributes significantly to the model’s camera pose control.
@@ IV-C Ablation Study
Video Generation Quality. We evaluate how the removal of each loss affects the realism and consistency of the generated video. This includes assessing structural coherence, temporal smoothness, and fidelity to the intended viewpoint. The quality of video generation is measured using the Fréchet Inception Distance (FID) [39].
@@ IV-D Cross-domain Application
In previous experiments, we demonstrate that the introduced PosePilot significantly improves pose controllability in world models for autonomous driving. Surprisingly, when we apply this module to the general video generation task, it further enhances the precision of camera pose control, achieving a new state-of-the-art performance on RealEstate10K [38], as shown in Table III. This finding indicates that our approach possesses strong cross-domain adaptability and scalability, paving the way for more advanced video generation tasks in diverse scenarios.
@@ V Conclusion
In this paper, we introduce PosePilot, a lightweight yet powerful framework that significantly enhances camera pose controllability in generative world models. Drawing inspiration from self-supervised depth estimation, PosePilot leverages structure-from-motion principles to establish a tight coupling between camera pose and video generation. We incorporate self-supervised depth and pose readouts, allowing the model to infer depth and relative camera motion directly from video sequences. These outputs drive pose-aware frame warping, guided by a photometric warping loss that enforces geometric consistency across synthesized frames. Extensive experiments on autonomous driving and general-domain video datasets demonstrate that PosePilot significantly enhances structural understanding and motion reasoning in both diffusion-based and auto-regressive world models. By steering camera pose with self-supervised depth, PosePilot sets a new benchmark for pose controllability, enabling physically consistent, reliable viewpoint synthesis in world models.

### v2

@@ Abstract
world model, controllable video generation, autonomous driving
@@ I Introduction
In recent years, great advancements have been made in the development of autonomous driving systems. While existing AD methods [1, 2, 3, 4, 5, 6, 7, 8, 9] demonstrate promising results in various scenarios, they still face significant challenges when confronted with long-tail distribution or out-of-distribution situations. These edge cases, which occur infrequently in typical training datasets but are critical in real-world driving environments, expose limitations in the robustness and generalization ability of current models. One promising solution is world models [10, 11, 12, 13, 14, 15, 16, 17], which capture the structure and dynamics of an environment, enabling an autonomous driving system to simulate and predict future states by ”imagining” the external world. Through building such models, these generative models can help AD systems anticipate future states, reason about complex dynamics and better generalize to novel or unexpected situations.
@@
  • We propose PosePilot, a lightweight, plug-and-play module that injects camera pose controllability into generative world models. It supports flexible viewpoint control across both diffusion and autoregressive architectures.
  • We design a photometric warping supervision strategy that leverages self-supervised depth and ego-motion estimation to enforce forward and inverse geometric consistency in video generation.
  • Extensive experiments and ablations on driving and real-world datasets show that PosePilot improves pose accuracy, structural coherence, and generalization across diverse generative models.
@@ II-A World Models for Driving Scenes
Generative world models World Models [27, 28, 29, 30] have emerged as a promising paradigm for simulating complex driving environments by predicting future visual states. Recent advances enabling video generation to augment autonomous driving data with diverse road conditions and rare corner cases [31, 32, 33]. Existing methods use various conditions including images, text, 3D layouts, and actions to generate specific driving scenes or predict future video sequences. GAIA-1 [34] generates realistic and diverse driving videos by integrating video, text, and action inputs, demonstrating a strong understanding of contextual information and physical principles. DriveDreamer [35] improves driving scenario generation by adding HD maps and 3D boxes for better video quality and enabling the generation of future driving actions, aiding in decision-making. WorldDreamer [36] frames world modeling as an unsupervised challenge within visual sequence modeling, inspired by large language models. MUVO [37] enhances world modeling by integrating LIDAR point clouds, improving the prediction of driving environments and generating 3D occupancy grids that integrate well with downstream tasks. OccWorld [38] leverages 3D occupancy data to predict environmental evolution and guide autonomous vehicle actions, marking a shift toward multimodal approaches in autonomous driving research. In contrast, our work does not design a brand new framework for world model but serves as a plug-and-play module that builds upon existing world models to provide effective camera pose control during video generation for driving scenes.
@@ III Methodology
We propose PosePilot, a lightweight and modular framework designed to enhance pose controllability in generative world models. The core idea as shown in Fig. 2 is to inject self-supervised geometric supervision into the video generation process, enabling the model to reason about camera motion in a structurally grounded way. We adapt PosePilot on top of the powerful generative world models [44, 17, 45].
III-A Pose-aware Warping via Self-supervised Geometry
In contrast to traditional video generation models, a key factor in the world models for autonomous driving is the camera parameters. Once the extrinsic matrices (or pose) of each camera are avaiGiven two framesle, we can use the structure-from-motion methods to generate the structural and dynamic context to guide the video generation. Inspired by monocular self-supervised depth estimation [24], we incorporate a photometric control loss into our generative model. The objective of the photometric control loss is to derive real-world scale information from the camera’s extrinsic matrices, which provides prior information for video generation.
@@ III-A Pose-aware Warping via Self-supervised Geometry
III-B Bidirectional Consistency for Inverse Warping
Forward-only supervision may suffer from occlusions and partial observability. To reinforce geometric alignment, we introduce inverse photometric warping, which enforces consistency from generated frames back to the original viewpoints. This bidirectional alignment helps reduce the artifacts due to occlusions and dis-occlusions.
We denote the generated frames as $(\hat{f}^{i},\hat{f}^{j})$, which are obtained from the world model. The learnable depth readout provides their respective depth maps $(D^{i},D^{j})$, and the pose readout outputs the relative 6D camera transformations $T^{i\rightarrow j}$ and $T^{j\rightarrow i}$. To map $\hat{f}^{j}$ back to $\hat{f}^{i}$, we warp usxing the same bilinear sampling mechanism as above. Specifically, the pixel projection from $\hat{f}^{j}$ to the warped inverse view $\hat{f}^{j\rightarrow i}$ follows the same procedure as in the forward warping. Denoting $x^{j}$ as a pixel in the generated frame $\hat{f}^{j}$, its inverse-warped coordinate $x^{j\rightarrow i}$ is found by:
@@ III-B Bidirectional Consistency for Inverse Warping
III-C Refining Pose Estimates with Regression Supervision
To further improve pose precision, we apply direct supervision to the pose readout module using a mean squared error (MSE) loss. Given a reference pose $T_{ref}$, the predicted pose $\hat{T}$ is refined via:
$L_{\text{MSE}}=\left\|\hat{T}-T_{\text{ref}}\right\|_{2}^{2}$ (5)
This loss reduces drift in relative pose estimation and provides stable gradients for training under large motion variation.
III-D Overall Objective
We combine the generative loss $L_{g}$ (e.g., diffusion or autoregressive objective) with the three geometric consistency terms into the final objective:
@@ III-D Overall Objective
Here, $\alpha_{p}$, $\alpha_{p^{\text{inv}}}$ and $\alpha_{mse}$ are weights that balance fidelity and geometric alignment. In practice, these terms guide the generator to synthesize high-quality, structurally plausible videos that accurately reflect the specified camera trajectory.
@@ IV-B Main Results
We adapt our methods to the existing world model solutions on the nuScenes validation set. We divide the nuScenes validation set into four subsets. The TranErr and RotErr scores are measured on each subset and then averaged. We conduct experiments on most of the publicly available world models, including diffusion-based methods [44, 17] and auto-regression-based approach [45]. The results are shown in Tab. I. We can see that with our PosePilot, all of the models can get relatively lower TranErr & RotErr, which demonstrates the effectiveness of the proposed architecture. For example, in diffusion-based methods, with our PosePilot, Vista [17] gets a 0.31 decrease on translation error and 0.21 decrease on rotation error. In auto-regression-based approach, Drivingworld [45] also gets a 0.22 decrease on translation error and 0.16 decrease on rotation error, indicating the superior generalization of our method. We also show the qualitative results of PosePilot at Fig. 3, we can see that our PosePilot  contributes significantly to the model’s camera pose control.
@@ IV-C Ablation Study
Video Generation Quality. We evaluate how the removal of each loss affects the realism and consistency of the generated video. This includes assessing structural coherence, temporal smoothness, and fidelity to the intended viewpoint. The quality of video generation is measured using the Fréchet Inception Distance (FID) [49]. We find that removing either photometric loss leads to a noticeable increase in FID, indicating degraded visual fidelity and weakened viewpoint alignment. In particular, the inverse loss contributes to improved temporal consistency, reducing flickering and motion discontinuities across frames. This highlights the role of geometric supervision not only in pose accuracy, but also in maintaining coherent, scene-consistent synthesis.
@@ IV-D Cross-domain Application
In previous experiments, we demonstrate that the introduced PosePilot significantly improves pose controllability in world models for autonomous driving. Surprisingly, when we apply this module to the general video generation task, it further enhances the precision of camera pose control, achieving a new state-of-the-art performance on RealEstate10K [48], as shown in Table III. This finding indicates that our approach possesses strong cross-domain adaptability and scalability, paving the way for more advanced video generation tasks in diverse scenarios. Its reliance on self-supervised geometric signals makes it inherently adaptable across domains without task-specific tuning.
@@ V Conclusion
In this paper, we presented PosePilot, a lightweight and generalizable framework that significantly improves camera pose controllability in generative world models. By leveraging self-supervised structure-from-motion cues, PosePilot introduces a principled mechanism to couple depth and ego-motion estimation with video generation. Our approach enforces geometric consistency through pose-aware and inverse warping losses, while direct pose supervision further enhances viewpoint precision. Extensive experiments across autonomous driving and general-domain datasets show that PosePilot improves structural fidelity and motion alignment in diverse architectures. As a plug-and-play module, PosePilot requires no architectural changes and generalizes well across domains. We believe it offers a new perspective on integrating geometric priors into generative modeling, paving the way for controllable, physically grounded video synthesis.